### v1

# Fairness and Bias in Multimodal : A Survey

Abstract
The importance of addressing fairness and bias in artificial intelligence (AI)systems cannot be over-emphasized. Mainstream media has been awashed with news of incidents around stereotypes and bias in many of these systems in recent years. In this survey, we fill a gap with regards to the minimal study of fairness and bias in Large Multimodal Models (LMM s) compared to Large Language Models (LLM s), providing 50 examples of datasets and models along with the challenges affecting them; we identify a new category of quantifying bias (preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic; we critically discuss the various ways researchers are addressing these challenges. Our method involved two slightly different search queries on Google Scholar, which revealed that 33,400 and 538,000 links are the results for the terms "Fairness and bias in Large Multimodal Models" and "Fairness and bias in Large Language Models", respectively. We believe this work contributes to filling this gap and providing insight to researchers and other stakeholders on ways to address the challenge of fairness and bias in multimodal AI.
1 Introduction
Fairness and bias are very important topics that cut across many domains in the society. The rapid advancements in the research and applications of artificial intelligence (AI)have made them even more compelling in recent times, such that many studies have emerged on them Frankel and Vendrow (2020); Booth et al. (2021); Adewumi et al. (2022); Teo et al. (2024). One important gap in the literature, however, is that there is minimal study or survey on "Fairness and bias in Large Multimodal Models." By multimodal AI, we mean the dataset or AI model that can take one or more modalities as input and/or another as output. This gap is evidenced by the fact that there is a dearth of work on the topic. For example, an exact-phrase search on Google Scholar returns 33,400 links compared to 538,000 links for "Fairness and bias in Large Language Models." [^1] June 10, 2024 This implies more than 16 times the result for the former. Filtering the publication year range to 2014-2024 reduces the links to 17,700 and 20,300, respectively. We intend to contribute in filling that gap in this work.
The two terms, fairness and bias, are strongly related but fairness is concerned with equality and justice while bias is concerned with systematic error, which may arise from human prejudices Booth et al. (2021); Alkhaled et al. (2023). For the purpose of this survey, fairness may be defined as equal representation with regards to a given Sensitive Attribute (SA)Hutchinson and Mitchell (2019); Frankel and Vendrow (2020). Hence, we may consider a generative artifical intelligence (GenAI)to be fair if it generates both male and female samples with equal probabilities, with regards to the SA gender Teo et al. (2024). Bias is a (non-random) systematic error in a measurement resulting in a difference in accuracy in one group compared to another, given the ground truth Booth et al. (2021); Scheuneman (1979). We acknowledge there are other quantitative definitions of fairness and bias, as noted by Hutchinson and Mitchell (2019) and Weidinger et al. (2022).
It appears the emergence of big data, which has brought rapid advancement in the state-of-the-art (SotA), also brought along the increase in poor quality content and prediction, such as the increased criminal prediction for Black and Latino people observed by Birhane et al. (2024a). Similar issues are observed across many domains, including healthcare, employment, forensics, criminal justice, credit scoring, and computational social science, among others Liang et al. (2021); Ferrara (2023); Landers and Behrend (2023); Han (2023). According to Wolfe et al. (2023), the model VQGAN-CLIP, similarly to Stable Diffusion, generated sexualized images for the harmless prompt "a 17 year old girl," 73% of the time. The comparison to a similar prompt with the term "girl" replaced with "boy" shows a sharp contrast.
In view of the foregoing gap and challenges, this work critically surveys the literature with the primary objective of ascertaining what the state of work is on fairness and bias in multimodal AI, thereby making the following key contributions.
  1. 1.
    We fill the gap with a comprehensive survey of fairness and bias across a wide spectrum of LMM s, LLM s, and multimodal datasets, providing 50 examples of datasets and models in a structured way, along with the challenges affecting them.
  2. 2.
    We identify a new category of quantifying bias (i.e. preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic Ramesh et al. (2023); Cabello et al. (2023).
  3. 3.
    We critically discuss many important approaches to addressing the challenges of fairness and bias.
The rest of this paper is organised as follows. In the following Section (2), we discuss fairness and bias in AI, highlighting some of the theories. In Section 3, we explain the method used for this survey. In Section 4, we focus on various works along the two paradigms of LMM s and LLM s, discussing the datasets and models in the literature. In Section 5, we discuss widely on the methods to evaluate fairness and bias, the datasets for evaluation, and the debiasing strategies. We conclude the survey in Section 6 with a summary and possible future work.
2 Fairness and Bias in AI
2.1 Concepts of Fairness and Bias
A discussion about every possible type of fairness or bias is beyond the scope of this study, however, we refer readers to Mehrabi et al. (2021) for some of the discussion. In this section, we highlight some of the concepts around fairness and bias found in the Social Science literature.
2.1.1 Justice Theory
The theory is regarded as a three-part framework of distributive, interactional, and procedural justice perceptions Greenberg (1990); Landers and Behrend (2023). Distributive justice, when outcomes are expected to be distributed equally, is the overarching aspect related to AI fairness and bias, according to Landers and Behrend (2023). It is based on equality rules, need, or equity, which are influenced by social and cultural values.
2.1.2 Equity Theory
Equity theory uses a unidimensional concept of fairness instead of multidimensional, according to Leventhal (1980). It perceives justice solely on the merit principle and the final distribution of reward (or punishment), where reward is proportional to contribution Adams and Freedman (1976).
2.1.3 Objectification Theory
Just as inanimate objects have no emotions or thoughts, objectification theory establishes a view of a subject as primarily without human characteristics, especially for women and girls Fredrickson and Roberts (1997); Heflick et al. (2011); Andrighetto et al. (2019). The theory identifies sexual objectification bias, which is when the emotions or thoughts of a person are disregarded and one is treated as mere body parts for sex Fredrickson and Roberts (1997); Wolfe et al. (2023).
2.2 Consequences & Legal Perspectives
Fredrickson and Roberts (1997) confirms that objectification victimises the subject and may result in habitual body monitoring, thereby increasing mental health risks, sexual dysfunction, eating disorders, and depression. This unhealthy reality is also confirmed by Swim et al. (2001). They realized that sexist incidents occur more against women and have negative emotional consequences for them. Some of these incidents are traditional gender role stereotypes, degrading remarks, and sexual objectification, which are found in the data used for training AI models. For details about the mechanics of training language models, we refer readers to Radford et al. (2019) and Hoffmann et al. (2022). It is not surprising, therefore, that the use of these models cause the same negative effects for those affected. Hence, fairness and bias are not only ethical or moral issues but have legal implications Landers and Behrend (2023). In the United States (US), disparate treatment because of sensitive attributes is unlawful Berry (2015); Hutchinson and Mitchell (2019); Meng et al. (2022). This is also the case in some other countries Zafar et al. (2017).
3 Method
To address the gap identified earlier, our method involved the simple approach of filtered search on Google scholar with two slightly differently-worded phrases: (1) "Fairness and Bias in Large Multimodal Models" and (2) "Fairness and Bias in Large Language Models" and they returned 17,700 and 20,300 links,1 respectively, over many pages. The search was filtered to the period 2014-2024 during which deep learning made significant progress. Apparently, there are fewer scientific papers on the former. We compare with "Language Models" only because these, with the introduction of the Transformer architecture Vaswani et al. (2017), have influenced computer vision Yuan et al. (2021) and serve as core component for many multimodal AI.
We captured archived papers (e.g. arXiv) but then thoroughly reviewed pair-reviewed papers that were returned on the first 10 pages (100 links) for each search text and identified the publishers (or venues), as listed in Table 1. We note that the more relevant papers turn up on the first page of the search while the quality of results degrade as one progresses through the pages. From the Table, it can be observed that there are 36 pair-reviewed papers on multimodal compared to 39 on language, among the 200 papers. We note that Google Scholar sometimes returns the arXiv version of a paper instead of the peer-reviewed version, hence, it’s very useful to include the data from arXiv. In total, 55 papers were published under multimodal compared to 88 for language. Each paper was reviewed for their contributions, including the datasets, models, possible solutions proffered on fairness and bias, and other relevant discussions.
Table 1: Distribution of scientific papers (2014-2024)
# Publisher Multimodal Language
1 IEEE 11 2
2 ELSEVIER 2 2
3 ACM 9 9
4 Springer 1 1
5 NeurIPS 4 8
6 Nature 3 1
7 MDPI - 1
8 MLR 3 5
9 PKP 2 -
10 MIT Press - 4
11 PubMed 1 -
12 Cambridge - 1
13 Patterns - 1
14 PLOS - 1
15 JMLR - 2
16 Science - 1
Sub-total 36 39
17 arXiv 19 49
Total 55 88
4 Fairness and Bias in LLM s and LMM s
In this section we review some LMM s and LLM s and the fairness and bias challenges they have. Tables 2 and 3 summarize some relevant datasets and the models, respectively. All the 25 datasets identified have their challenges and by extension the 25 AI models they are trained on. Some of these challenges include stereotypes, porn, misogyny, racial, gender, religious, cultural, age, and demographic biases. It is commonly agreed AI models learn much of their bias from the data they are trained on and many datasets, especially those for pretraining, are from the Internet, which contains a diverse spectrum of content Wolfe and Caliskan (2022).
4.1 LMM s
Table 2: Summary of Some Datasets and Their Fairness & Bias Challenges (Data in the second part of the table are usually used in downstream tasks).
# Dataset Modality Some Challenges
1 CommonCrawl Raffel et al. (2020) Text & Vision Fake news, hate speech, porn & racism Gehman et al. (2020); Luccioni and Viviano (2021)
2 LAION-400M & 5B Schuhmann et al. (2021, 2022) Text & Vision Misogyny, stereotypes & porn Birhane et al. (2021, 2024b)
3 WebImageText (WIT)Radford et al. (2021) Text & Vision Racial, gender biases Radford et al. (2021)
4 DataComp Gadre et al. (2024) Text & Vision Racial bias Gadre et al. (2024)
5 WebLI Chen et al. (2022) Text & Vision Age, racial, gender biases & stereotypes Chen et al. (2022)
6 CC3M-35L Thapliyal et al. (2022) Text & Vision Cultural bias Thapliyal et al. (2022)
7 COCO-35L Thapliyal et al. (2022) Text & Vision Cultural bias Thapliyal et al. (2022)
8 WIT Srinivasan et al. (2021) Text & Vision Cultural bias Srinivasan et al. (2021)
9 Colossal Cleaned CommonCrawl (C4) Raffel et al. (2020) Text Offensive language, racial bias Raffel et al. (2020)
10 The Pile Gao et al. (2020) Text Religious, racial, gender biases Gao et al. (2020)
11 CCAligned El-Kishky et al. (2020) Text Porn, racial bias El-Kishky et al. (2020)
12 OpenAI WebText Radford et al. (2019) Text Gender, racial biases Gehman et al. (2020)
13 OpenWebText Corpus (OWTC) Text Gender, racial biases Gehman et al. (2020)
14 ROOTS Laurençon et al. (2022) Text Cultural bias Laurençon et al. (2022)
15 VoxCeleb 1 Nagrani et al. (2020) Audio & Vision Demographic, gender biases Chung et al. (2018)
16 VoxCeleb 2 Chung et al. (2018) Audio & Vision Demographic, gender biases Chung et al. (2018)
17 First Impressions Escalante et al. (2020) Audio & Vision Racial, gender biases Yan et al. (2020)
18 XM3600 Thapliyal et al. (2022) Text & Vision Cultural bias Thapliyal et al. (2022)
19 VQA Antol et al. (2015) Text & Vision Gender bias Ruggeri and Nozza (2023)
20 VQA 2 Goyal et al. (2017) Text & Vision Gender bias Ruggeri and Nozza (2023)
21 MS COCO Lin et al. (2014) Text & Vision Gender bias Cabello et al. (2023); Zhao et al. (2017)
22 Multi30K Elliott et al. (2016) Text & Vision Racial bias Wang et al. (2022)
23 MIMIC-IV Johnson et al. (2023) Text & Vision Ethnic, racial, marital status biases Meng et al. (2022)
24 MAB Alkhaled et al. (2023) Text Racism, misogyny, stereotypes Alkhaled et al. (2023); Pagliai et al. (2024)
25 Twitter corpus Huang et al. (2020) Text Age, gender, racial biases Huang et al. (2020)
Wolfe et al. (2023) found evidence of sexual objectification bias in models based on Contrastive Language-Image Pretraining (CLIP). The 9 CLIP models that were investigated were trained on internet-wide web crawls. CLIP is known to be quite accurate Radford et al. (2021), however, it also appears to have scaled the biases inherent in its training data. Also, Wolfe and Caliskan (2022) found that more than Latino, Asian or Black, White persons are more associated with collective in-group words in embeddings from CLIP Radford et al. (2021), SLIP Mu et al. (2022), and BLIP Li et al. (2022), as measured with EATs. For a definitive assessment, their work would have benefited from additional experiments involving data of people outside the United States (US), since they used Chicago Face Database (CFD)Ma et al. (2015). CFD is a dataset of 597 people recruited in the U.S. Similarly, Teo et al. (2024) found that Stable Diffusion exhibits gender bias on slight changes to the prompts.
Besides work on text and text-visual multimodal systems or data, there are some work on audio-visual systems or data Fenu and Marras (2022). In the work by Fenu and Marras (2022), they perform comparative analysis on audio-visual speaker recognition systems, using fusion at the model step. They found that the highest accuracy and the lowest disparity across groups are achieved compared to unimodal systems.
In other works, Peña et al. (2023) evaluated AI-based recruitment for multimodal data but in a fictitious case study, which may limit its generalizability in real-world applications. Booth et al. (2021) performed a case study of automated video interviews and found that combining more than one modality increases bias and reduces fairness, similarly to what happens when scaling crawled data for training models. Cabello et al. (2023) investigated fairness and bias in LMM s. Other researchers investigated the impact of multimodal data/models on personlity assessment Yan et al. (2020), cyberbullying Alasadi et al. (2020), health records Meng et al. (2022) and more Birhane et al. (2021); Ferrara (2023); Zhang et al. (2022)
4.2 LLM s
Table 3: Summary of Models and Some of Their Fairness & Bias Challenges. (*modified for audio-visual)
# LMM s Modality Training Data Some Challenges
1 VQGAN-CLIP Crowson et al. (2022) Text & Vision WIT & ImageNet Misogyny, racial bias Pagliai et al. (2024); Wolfe and Caliskan (2022)
2 DALL-E 2 Ramesh et al. (2021) Text & Vision Conceptual Captions Occupational stereotypes, gender bias Ramesh et al. (2022); Mandal et al. (2023)
3 GLIDE Nichol et al. (2022) Text & Vision WIT & Conceptual Captions Gender stereotypes Nichol et al. (2022)
4 Stable Diffusion Rombach et al. (2022) Text & Vision LAION-5B Gender bias Teo et al. (2024)
5 SLIP Mu et al. (2022) Text & Vision YFCC100M Racial bias Wolfe and Caliskan (2022)
6 CLIP Radford et al. (2021) Text & Vision WIT Racial bias Radford et al. (2021); Wolfe and Caliskan (2022)
7 BLIP Li et al. (2022) Text & Vision MS COCO, Conceptual Captions & LAION-400M Racial, gender & age biases Wolfe and Caliskan (2022); Ruggeri and Nozza (2023)
8 PaliGemma Text & Vision WebLI, CC3M-35L , WIT Porn, offensive language222ai.google.dev/gemma/docs/paligemma/model-card
9 PaLI-3 Chen et al. (2022) Text & Vision WebLI Age, racial, gender biases & stereotypes Chen et al. (2022)
10 Falcon 2 Almazrouei et al. (2023) Text & Vision RefinedWeb Harmful content, cultural bias Almazrouei et al. (2023)
11 BEiT Wang et al. (2023) Text & Vision Conceptual 12M, ImageNet-21K, Wikipedia Gender, cultural biases Brinkmann et al. (2023)
12 LLaVA Liu et al. (2024b, a) Text & Vision Conceptual Captions Cultural bias Liu et al. (2024b)
13 ResNet-50* He et al. (2016) Audio & Vision ImageNet Gender bias Fenu and Marras (2022)
14 GPT4o Achiam et al. (2023) Text, Audio & Vision WebText, Github, etc Stereotypes, racial bias Aich et al. (2024)
LLM s
15 GPT3 Brown et al. (2020) Text CommonCrawl & WebText Gender, racial, religious biases Brown et al. (2020); Gehman et al. (2020)
16 PaLM Chowdhery et al. (2024) Text Wikipedia, social media, Github Occupation, gender, sexual, religious biases Chowdhery et al. (2024)
17 LaMDA Thoppilan et al. (2022) Text Social media & Wikipedia Gender bias Thoppilan et al. (2022)
18 GLaM Du et al. (2022) Text Wikipedia & social media Toxicity, gender bias Du et al. (2022)
19 GPT2 Radford et al. (2019) Text WebText Sexual, racial, gender biases Sheng et al. (2019); Gehman et al. (2020)
20 LLaMA-3 Text web text & Github Stereotypes, gender, racial, sexual, religious, biasesAich et al. (2024); Touvron et al. (2023)
21 LLaMA-2 Touvron et al. (2023) Text web text Toxicity, gender, racial, sexual, religious, biases Touvron et al. (2023)
22 CTRL Keskar et al. (2019) Text Wikipedia, Project Gutenberg, OpenWebText Gender, racial biases Gehman et al. (2020)
23 Aurora-M Nakamura et al. (2024) Text The Pile Offensive language, religious, racial, gender biases Gao et al. (2020); Nakamura et al. (2024)
24 Mixtral-8x7B Jiang et al. (2024) Text web text Stereotypes, racial, gender, occupational biases Jiang et al. (2024); Aich et al. (2024)
25 BLOOM Le Scao et al. (2023) Text ROOTS Toxicity, gender, religious, disability, age biases Le Scao et al. (2023)
The introduction of the Generative Pretrained Transformer (GPT-2) Radford et al. (2019) was a turning point in the LLM lanscape with its 1.5B parameters. As pointed out earlier, training such a large model required a lot of data and the Internet-sourced WebText was used for this purpose. Updated versions of the dataset have also been used for its recent successors, including GPT-3 Brown et al. (2020) and GPT-4 Achiam et al. (2023). The attendant problems of bias witnessed in GPT-2 Gehman et al. (2020) have followed successive versions.
Quantitative bias measurement in NLP may be placed into 3 categories: preuse, as carried out by Alkhaled et al. (2023) and Adewumi et al. (2023b), intrinsic Caliskan et al. (2017); May et al. (2019), and extrinsic, as observed by Ramesh et al. (2023). The first, second and third involve quantifying bias in the training data, in the trained model’s representation, and in the outputs of the downstream task of the model, respectively. Far more work has focused on the latter two than the first and gender bias than other dimensions Ramesh et al. (2023). For example, Delobelle et al. (2022) and Welbl et al. (2021) measured bias in pre-trained language models. Others have surveyed fairness and bias in LLM s Bender et al. (2021); Meade et al. (2022); Gallegos et al. (2023); Chang et al. (2024); Myers et al. (2024).
In other works, Schramowski et al. (2022) show that moral directions, i.e. what is morally right or wrong to do, are present in LLM s. It is, however, highly debatable if models can be considered moral agents. Also, Santurkar et al. (2023) explored whose opinions LLM s reflect. With regards to healthcare and bias, work has also been done Harrer (2023); Nashwan and Abujaber (2023). In a collaborative effort, bench authors (2023) probed LLM s in the BIG-Bench of 200 tasks, including many that are related to bias.
Bias in multilingual AI:
Some of the multimodal data in Table 2 contain multilingual data. Such result in multilingual models and embeddings. For example, CC3M-35L, COCO-35L, and WebLI, used to train PaLI. WebLI is a mix of pre-training tasks with texts in 109 languages Chen et al. (2022). Some of these languages fall in the category of low-resource languages Adewumi et al. (2023a). Kurpicz-Briki (2020) reports statistically significant bias in German word embeddings based on the origin of a name in relation to pleasant and unpleasant words, using WEAT Caliskan et al. (2017). In the study by Wambsganss et al. (2022), they found that the pretrained German language models, GermanBERT, GermanT5, and German GPT-2, had substantial conceptual, racial, and gender bias. This was also confirmed by Kraft et al. (2022), who observed sexist stereotypes in some of the models (e.g. family- and care-related terms related to female while crime and perpetrators related to male). Similarly, for Dutch, Delobelle et al. (2020) investigated gender and occupation biases in RobBERT (a Dutch RoBERTa Liu et al. (2019)), through a template-based association test Kurita et al. (2019); May et al. (2019). Huang et al. (2020) also identified biases related to people’s origin and age in Italian, English, Polish, Portuguese, and Spanish, using a Twitter corpus.
5 Discussion
Although there are many limitations or risks of multimodal AI or LLM s Acosta et al. (2022); Adewumi et al. (2024b); Pettersson et al. (2024); Adewumi et al. (2024a), perhaps, the issue of fairness and bias rank among the topmost Mehrabi et al. (2021). In addition, in the taxonomy of 21 risks of language models provided by Weidinger et al. (2022), the first category is "Discrimination, Hate speech and Exclusion". Given, the recurring challenges in this regard, we are of the view that existing tools for evaluating or handling fairness and bias in these systems need to be improved Zhao et al. (2018); Rudinger et al. (2018). It may be almost impossible to automatically filter a dataset or debias a model to be 100% free of unfair, bias or toxic content but the research community and other stakeholders may need to determine what levels are acceptable and if it should be a requirement to have human-in-the-loop methods. In this section, we discuss further methods to audit or evaluate fairness and bias, datasets for such evaluation, and debiasing strategies. We hope that such discussion will spur more researchers and stakeholders to see the critical importance of AI that is fair and free, as much as possible, from bias.
5.1 Methods to audit, measure, and evaluate fairness and bias
Embedding Association Tests (EATs)have been used in several studies Caliskan et al. (2017); Kurpicz-Briki (2020); Wolfe et al. (2023), starting with the introduction of Word Embedding Association Test (WEAT)by Caliskan et al. (2017) and improvements in Sentence Embedding Association Test (SEAT)May et al. (2019) and Relational Inner Product Association (RIPA)Ethayarajh et al. (2019). Despite its widespread use, WEAT has the disadvantage that it may systematically overestimate the bias in a model. Another embedding evaluation method is cosine similarity. It was used for zero-shot classification by Radford et al. (2021). It may also be used to audit fairness and bias by evaluating the similarity in image and text embeddings Wolfe et al. (2023). The visual tool, Gradient-weighted Class Activation Mapping (GRAD-CAM), generates a saliency map, which shows the most relevant regions of an image for given attributes Selvaraju et al. (2017); Wolfe et al. (2023). In an evaluation carried out by Wolfe et al. (2023), they discovered that the computed average saliency maps included only face regions for non-objectified images but both face and chest regions for objectified images, in a possible indication of sexual objectification bias. Not Safe For Work (NSFW)detector uses a tag alongside each image for filtering undesirable content Schuhmann et al. (2021); Birhane et al. (2024b)
A recent metric introduced by Alkhaled et al. (2023) is bipol. It uses a two-step procedure in estimating bias in data Adewumi et al. (2023b); Pagliai et al. (2024). Bipol has the weakness that if the bias classifier is not accurate enough, false positives will weaken the evaluation score. Another measure is Area Under the Curve (AUC), as used by Meng et al. (2022) in the investigation of algorithmic fairness of mortality prediction, where they noted that Machine Learning (ML)methods obtain lower scores, usually, when it involves groups with higher mortality rates. Teo et al. (2024) proposed CLassifier Error-Aware Measurement (CLEAM), a framework for better performance in bias estimation, while Booth et al. (2021) measured gender bias using accuracy of Spearman rank-based correlation ($\rho$). Furthermore, Nozza et al. (2021) introduced the score “HONEST”, which was tested with respect to gender bias in text generation in 6 languages: Italian, French, Portuguese, Romanian, Spanish and English. It measures the probability that a language model will output hurtful text given a certain template and lexicon.
5.2 Datasets for Bias Evaluation
Different datasets have been introduced for bias evaluation. Esiobu et al. (2023) introduced 2 novel datasets AdvPromptSet and HolisticBiasR, with which they evaluated 12 demographic dimensions for different LLM s. Ruzzante et al. (2022) introduced Sexual OBjectification and EMotion Database (SOBEM)for sexual objectification bias studies. It consists of 280 pictures of objectified and non-objectified female models with 3 different emotions and a neutral face. Bias Benchmark for QA (BBQ)is a question-sets dataset that employs templates crafted to reflect specific biases identified in society. It was introduced by Parrish et al. (2022) and aims to expose implicit prejudices that may exist against individuals from legally protected categories. BEAVERTAILS was introduced by Ji et al. (2024). It assesses question-answer pairs, tested on LLM s, with regards to 14 different harm categories, which are not exclusive. Discrimination, Stereotype, Injustice make up the second category in the list. Additional datasets for bias evaluation include RedditBias by Barikeri et al. (2021), RealToxicityPrompts, which comprises of 100K English sentences Gehman et al. (2020), HarmfulQ for zero-shot Chain of Thought (CoT)across stereotype benchmarks and harmful questions Shaikh et al. (2023), and Bold Dhamala et al. (2021).
5.3 Debiasing strategies on data, algorithms, procedures
5.3.1 Curate over Crawl
One important method to address bias in datasets will be to curate over crawl. This is especially so because web crawling has been the popular approach to getting internet data in the shortest possible time Birhane et al. (2021). The assumption of scale beats noise is the rationale for this approach by some researchers Jia et al. (2021). Unfortunately, this misconception about scaling does not only scale the quality part of the dataset but the noise along with it, no matter how small, as shown by Birhane et al. (2024b) when they observed 12% increase in hate content due to scaling. Clearly, despite the advantage of curation, one hurdle to overcome with the curate over crawl approach will be the issue of scaling.
5.3.2 Counterfactual Data Augmentation (CDA)
Zhao et al. (2018) used CDA to show that it removes bias with minimal performance degradation on coreference benchmarks when combined with existing word-embedding debiasing methods. It involves generating alternate examples of what exists (counterfactual) of data points to counter or mitigate bias. This method has gained attention in the field Meade et al. (2022); Barikeri et al. (2021).
5.3.3 Improved Filtering
It has been shown that poor filtering during the data creation process allows low quality data in the final dataset Birhane et al. (2024a). It may not be possible to automatically filter large data to be 100% fit for purpose but improving the existing methods of filtering will go a long way in mitigating bias.
5.3.4 Other strategies
Oversampling under-represented groups in the data has been advocated by Buolamwini and Gebru (2018). There are generation detoxifying methods with the potential to reduce bias Gehman et al. (2020). These include data-based further pretraining and decoding-based generation. Similarly, Cabello et al. (2023) showed that continued pretraining on gender-neutral data improves fairness by reducing group disparities in some language-vision tasks. In another work, Yan et al. (2020) used adversarial learning. They added a discriminator to jointly learn with the predictor for the sensitive attributes. Meanwhile, Zayed et al. (2024) addressed fairness by pruning in LLM s while Guo et al. (2022) introduced auto-bias, by directly probing the biases in pretrained models through prompts. Liang et al. (2021) introduced the Autoregressive Iterative Nullspace Projection (A-INLP)method to carry out post-hoc debiasing on LLM s. Additionally, there are Self-Debias Schick et al. (2021), Hard-Debias Bolukbasi et al. (2016), SentenceDebias Liang et al. (2020) and Dropout methods Webster et al. (2020); Meade et al. (2022)
6 Conclusion
Fairness and bias are very important considerations in multimodal AI. In this work, we presented the challenges of fairness and bias in multimodal data, LMM s, and LLM s, defining what both terms mean within the scope of this survey, while acknowledging other definitions in the literature. We discussed the concepts of fairness and bias from the perspective of the Social Science and showed the distribution of scientific publications across 17 publishers, which reveals the gap of study in large multimodal AI compared to LLM s, which this work contributes to filling. Our discussion around the methods to measure fairness and bias, datasets for evaluation, and debiasing strategies will provide researchers and other stakeholders with insight on how to approach these issues. For future work, it will be worthwhile to re-evaluate the progress made with the metrics and tools for quantifying and mitigating fairness and bias concerns, respectively.
Acknowledgements
This work is funded by the EU HumaneAI Project. The work is partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU).